# On Modular Invariants of Truncated Polynomial Rings in low ranks

### v1

@@ 3. A lower bound on the total dimension of the invariant subspaces
Note that all coefficient polynomials $g_{j}$ are $\operatorname{GL}_{2}$-invariant in $x_{2}$ and $x_{3}$. If $i(q-1)=q^{m}-1$, then $x_{1}^{q^{m}-1}g(x_{2},x_{3})$ is a linear combination of polynomials from the last 3 families of $\mathcal{B}_{m}(1,2)$. We can then subtract from $f$ this polynomial and can now assume that $i(q-1)\leq q^{m}-q$.
Since $f$ is a $B$-invariant, it is a linear combination of polynomials of the form $x_{1}^{j_{1}(q-1)}\delta_{2}(Q_{2,1}^{j_{2}})$ or $Q_{1,0}^{j_{1}}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$. By direct inspection, the lowest $x_{1}$-degree of these polynomials are $x_{1}^{j_{1}(q-1)}x_{2}^{q^{m}-q}x_{3}^{j_{2}q(q-1)}$ and $x_{1}^{j_{1}(q-1)}x_{2}^{j_{2}q(q-1)}Q_{2,1}^{qj_{3}}(x_{2},x_{3})$ respectively. In both cases, the exponent of $x_{2}$ and $x_{3}$ in the coefficient of the lowest $x_{1}$-degree is divisible by $q(q-1)$.
We conclude that $g_{i}(x_{2},x_{3})$ is a polynomial in which all exponents of $x_{2}$ and $x_{3}$ are divisible by $q(q-1)$. Since $g_{i}\in{\mathcal{Q}}_{m}^{G}(x_{2},x_{3})$, it must be a $q$-th power of a polynomial in ${\mathcal{Q}}_{m-1}(2)$ in the variables $x_{2},x_{3}$.
On the other hand, the lowest $x_{1}$-degree of the first 3 families of $\mathcal{B}_{m}(1,2)$ are
@@
    $x_{1}^{j_{1}(q-1)}[x_{2}^{q^{m-1}-1}x_{3}^{q^{m-1}-1}]^{q}$.
Hence $x_{1}^{i(q-1)}g_{i}(x_{2},x_{3})$ is the lowest $x_{1}$-degree of a polynomial in the span of $\mathcal{B}_{m}(1,2)$. So by subtracting from $f$ an appropriate linear combination of polynomials in $\mathcal{B}_{m}(1,2)$, we obtain another $P(1,2)$-invariant polynomial whose smallest $x_{1}$-degree is strictly greater than that of $f$. By induction, the required result follows. ∎
@@
We can also prove directly $\mathcal{B}_{m}(1,2)$ is linearly independent. In fact, the generators can be distinguished by looking at the lowest $x_{1}$-degrees and the corresponding coefficients as given in the table below.
$\mathcal{B}_{m}(1,2)$ $x_{1}$-degree Coefficient
@@
We need to prove that $\mathcal{B}_{m}(3)$ is a generating set for ${\mathcal{Q}}_{m}^{G}(3)$. The proof’s strategy is the same as in the rank $2$ case, but the details are much more involved. The proof is divided into three steps. Using the transfer, we first construct a slightly bigger generating set $\mathcal{B}^{\prime}$ of the required form without restricting the Dickson polynomials:
@@
and as a result, ${\mathcal{D}}_{2}$ can be replaced by the much smaller subset $\Delta^{m}_{2}$. Finally, we prove that when restricted to ${\mathcal{Q}}_{m}(3)$, the rank $3$ Dickson algebra belongs to the subspace spanned by $\mathcal{B}_{m}(3)$. This is proved by first showing that all the edge monomials of $\Delta^{m}_{3}$ are in this span. Thus the subspace spanned by $\mathcal{B}_{m}(3)$ is a ${\mathcal{D}}_{3}$-submodule of ${\mathcal{Q}}_{m}^{G}(3)$. Since it contains $\Delta^{m}_{3}$, it will contain the entire ${\mathcal{D}}_{3}$.
@@
The set $\mathcal{B}^{\prime}$ consisting of 4 families below is a generating set for ${\mathcal{Q}}_{m}^{G}(3)$.
@@
We use the relative transfer from $P(1,2)$ to $G$. Among the six families in $\mathcal{B}_{m}(1,2)$, the last family is already a $G$-invariant. So are $\delta_{3}(Q_{2,1}^{j_{2}})$ and $\delta_{2}^{2}(1)$. On the other hand, $\operatorname{Tr}(x_{1}^{j_{1}(q-1)})$ is clearly an element of ${\mathcal{D}}_{3}$ and is a multiple of $Q_{3,0}$. It is known that
$Q_{3,0}\delta_{3}(Q_{2,1}^{j_{2}})=Q_{3,0}\delta_{2}^{2}(1)=0\quad\text{in ${\mathcal{Q}}_{m}$}.$
It follows that the transfer of the second and third family in $\mathcal{B}_{m}(1,2)$ contribute $G$-invariants of the form $\delta_{3}(Q_{2,1}^{j_{2}})$, $j_{2}<[m-1]_{q}$ and $\delta_{2}^{2}(1)$. The transfer of the 4th family clearly belongs to ${\mathcal{D}}_{3}$, so it remains to consider the 5th family.
We claim that if $j<[m]_{q}$, then
$\operatorname{Tr}_{P(1,2)}^{G}(x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3}))=a_{m,3,j+1}.$
Indeed, consider $x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3})$ as a $B_{3}$-invariant and taking the transfer from $B_{3}$ to $P(2,1)$. In Section 4, we have already proved that
$\operatorname{Tr}_{B_{2}}^{\operatorname{GL}_{2}}(x_{1}^{q^{m}-1}x_{2}^{a(q-1)})=y_{a+1}(x_{1},x_{2}).$
Hence
$\operatorname{Tr}_{B_{3}}^{P(2,1)}(x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3}))=\sum_{a=j}^{(q^{m}-q)/(q-1)}y_{a+1}(x_{1},x_{2})x_{3}^{q^{m}-q+j(a-1)-a(q-1)}.$
The last sum is exactly $a_{m,3,j+1}=\delta_{2}^{2}(Q_{1,0}^{j+1})$, which is already a $\operatorname{GL}_{3}$-invariant. By the transitivity property of the transfer, we have the required claim. The proof is finished. ∎
The next step is to show that in the third family, one need not use the entire ${\mathcal{D}}_{2}$. The following technical lemma will be used repeatedly.
@@
Let $s,t$ be non-negative integers and $0\leq i<[m]_{q}$. Then the following equality holds in ${\mathcal{Q}}_{m}(3)$:
@@
Recall that $\delta_{s+1}(f)$ is a polynomial if $f$ is a polynomial (possibly with higher than $s$ variables), which is $\operatorname{GL}_{s}$-invariant. Hence the expression under consideration is a polynomial. If $t\geq 1$, then by definition, $\delta_{2}(V_{1}^{s(q-1)}V_{2}^{t(q-1)}\delta_{2;m}(Q_{1,0}^{i}))$ equals
@@
The first summand is clearly trivial in ${\mathcal{Q}}_{m}$ because the quotient involved is a polynomial. For the second summand, it is also trivial if $s>1$, and is $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-1)(q-1)}$ if $s=1$. If $s=0$, the result is
$-(t(q-1)-1)x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-2)(q-1)}.$
Now suppose $t=0$. Then we have
$\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}^{s(q-1)}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{2}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}^{s(q-1)}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{1}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{1},x_{3})}\big{]}\\ =\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{2}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{1}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\big{]}+\\ \frac{x_{1}^{q^{m}}x_{2}^{q^{m}}x_{3}^{i(q-1)}}{L_{2}(x_{1},x_{2})}\Bigg{[}\frac{x_{2}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{2}^{q-1}-x_{3}^{q-1}}-\frac{x_{1}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{1}^{q-1}-x_{3}^{q-1}}\Bigg{]}$
The first summand is $\delta_{2}^{2}(Q_{1,0}^{s+i})$. In ${\mathcal{Q}}_{m}(3)$, the second summand is trivial when $s\leq 1$ and equals $-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(s-2+i)(q-1)}$ if $s\geq 2$. ∎
We are now ready to show that
@@
The subspace of ${\mathcal{Q}}_{m}^{G}(3)$ spanned by the following three families of elements

### v2

@@ 3. A lower bound on the total dimension of the invariant subspaces
Note that all coefficient polynomials $g_{j}$ are $G_{2}$-invariant in $x_{2}$ and $x_{3}$. If $i(q-1)=q^{m}-1$, then $x_{1}^{q^{m}-1}g(x_{2},x_{3})$ is a linear combination of polynomials from the last three families of $\mathcal{B}_{m}(1,2)$. We can then subtract from $f$ this polynomial and can now assume that $i(q-1)\leq q^{m}-q$.
Since $f$ is a $B$-invariant, it is a linear combination of polynomials of the form $x_{1}^{j_{1}(q-1)}\delta_{2}(Q_{2,1}^{j_{2}})$ or $Q_{1,0}^{j_{1}}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$. Their lowest $x_{1}$-degree summands are $x_{1}^{j_{1}(q-1)}x_{2}^{q^{m}-q}x_{3}^{j_{2}q(q-1)}$ and $x_{1}^{j_{1}(q-1)}x_{2}^{j_{2}q(q-1)}Q_{2,1}^{qj_{3}}(x_{2},x_{3})$ respectively. In both cases, the exponents of $x_{2}$ and $x_{3}$ in these summands are divisible by $q(q-1)$. So the same is true for $g_{i}(x_{2},x_{3})$. Since $g_{i}\in{\mathcal{Q}}_{m}(x_{2},x_{3})^{G}$, it must be a $q$-th power of a polynomial in ${\mathcal{Q}}_{m-1}(2)$ in the variables $x_{2},x_{3}$.
On the other hand, the lowest $x_{1}$-degree summand of the first three families of $\mathcal{B}_{m}(1,2)$ are
@@
    $x_{1}^{j_{1}(q-1)}\big{(}x_{2}^{q^{m-1}-1}x_{3}^{q^{m-1}-1}\big{)}^{q}$.
Hence $x_{1}^{i(q-1)}g_{i}(x_{2},x_{3})$ is the lowest $x_{1}$-degree part of a polynomial in the span of $\mathcal{B}_{m}(1,2)$. By subtracting from $f$ an appropriate linear combination of polynomials in $\mathcal{B}_{m}(1,2)$, we obtain another $P(1,2)$-invariant polynomial whose lowest $x_{1}$-degree is strictly greater than that of $f$. By induction, the required result follows. ∎
@@
We can also prove directly that $\mathcal{B}_{m}(1,2)$ is linearly independent. In fact, the generators can be distinguished by looking at the lowest $x_{1}$-degrees and the corresponding coefficients as given in the table below.
$\mathcal{B}_{m}(1,2)$ Lowest $x_{1}$-degree Coefficient
@@
We need to prove that $\mathcal{B}_{m}(3)$ is a generating set for ${\mathcal{Q}}_{m}(3)^{G}$. Our strategy is the same as in the rank $2$ case, but the details are much more involved. The proof is divided into three steps. Using the transfer, we first construct a slightly bigger generating set $\mathcal{B}^{\prime}$ of the required form without restricting the Dickson polynomials:
@@
and as a result, ${\mathcal{D}}_{2}$ can be replaced by the much smaller subset $\Delta^{m}_{2}$. Finally, we prove that when restricted to ${\mathcal{Q}}_{m}(3)$, the rank $3$ Dickson algebra belongs to the subspace spanned by $\mathcal{B}_{m}(3)$. This is proved by first showing that all the edge monomials of $\Delta^{m}_{3}$ are in this span. Thus the subspace spanned by $\mathcal{B}_{m}(3)$ is a ${\mathcal{D}}_{3}$-submodule of ${\mathcal{Q}}_{m}(3)^{G}$. Since it contains $\Delta^{m}_{3}$, it will contain the entirety of ${\mathcal{D}}_{3}$.
@@
The set $\mathcal{B}^{\prime}$ consisting of 4 families below is a generating set for ${\mathcal{Q}}_{m}(3)^{G}$.
@@
A generating set for ${\mathcal{Q}}_{m}(3)^{G}$ can be obtained by taking the image of the relative transfer
$\operatorname{Tr}_{P(2,1)}^{G}\colon{\mathcal{Q}}_{m}^{P(1,2)}\longrightarrow{\mathcal{Q}}_{m}^{G}$
from the parabolic subgroup $P(1,2)$ to $G$, applied to the basis of ${\mathcal{Q}}_{m}^{P(1,2)}$ given in Proposition 1.8.
First, observe that $Q_{1,0}^{j_{1}}Q_{3,2}^{i_{1}}Q_{3,1}^{i_{2}}$ is a genuine $P(1,2)$-invariant, and hence its transfer is a genuine $G$-invariant, i.e., a Dickson polynomial.
Next, consider the second family. Since $\delta_{3}(Q_{2,1}^{j_{2}})$ is already a $G$-invariant in ${\mathcal{Q}}_{m}(3)$, we have
$\operatorname{Tr}_{P(2,1)}^{G}(Q_{1,0}^{j_{1}}\delta_{3}(Q_{2,1}^{j_{2}}))=(\operatorname{Tr}_{P(2,1)}^{G}Q_{1,0}^{j_{1}})\delta_{3}(Q_{2,1}^{j_{2}}).$
On the other hand, the transfer $\operatorname{Tr}_{P(2,1)}^{G}(Q_{1,0}^{j_{1}})$ is a genuine Dickson polynomial divisible by $Q_{3,0}$ unless $j_{1}=0$. Therefore, if $j_{1}$ is positive, the whole expression vanishes in ${\mathcal{Q}}_{m}(3)$:
$\operatorname{Tr}_{P(2,1)}^{G}(Q_{1,0}^{j_{1}}\delta_{3}(Q_{2,1}^{j_{2}}))=0\quad\text{in ${\mathcal{Q}}_{m}$}.$
A similar argument applies to the third family. Since $\delta_{2}^{2}(1)$ is already a $G$-invariant and is annihilated by $Q_{3,0}$, its transfer also vanishes. Hence from the second and third families in $\mathcal{B}_{m}(1,2)$, only elements $\delta_{3}(Q_{2,1}^{j_{2}})$ (for $j_{2}<[m-1]_{q}$) and $\delta_{2}^{2}(1)$ contribute nontrivially.
Now we turn to the fourth family. Recall that
$\delta_{1}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})=x_{1}^{q^{m}-1}Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3}).$
We claim that
$\operatorname{Tr}_{P(1,2)}^{G}(\delta_{1}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}))=\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}+1}).$
Indeed, by Proposition 2.3, the right-hand side is $G$-invariant. Moreover, we use the identity:
$V_{3}(x_{2},x_{1},x_{3})=V_{2}(V_{2}(x_{2},x_{1}),V_{2}(x_{2},x_{3}))=V_{2}(x_{2},x_{3})^{q}-V_{2}^{q-1}(x_{2},x_{1})V_{2}(x_{2},x_{3}),$
from which it follows that
$L_{2}^{q}(x_{2},x_{3})=x_{2}^{q}V_{2}(x_{2},x_{3})^{q}=x_{2}^{q}V_{3}(x_{2},x_{1},x_{3})+Q_{2,0}(x_{1},x_{2})L_{2}(x_{2},x_{3}).$
Applying the Laplace expansion along the last row of the numerator, we obtain:
$\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}+1})=Q_{2,0}\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})-\delta_{2}(Q_{1,0}Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}).$
Now consider $x_{1}^{q^{m}-1}Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})$ as a $B_{3}$-invariant and take the transfer from $B_{3}$ to $P(2,1)$. From Section 4, we have
$\operatorname{Tr}_{B_{2}}^{G_{2}}(x_{1}^{q^{m}-1}x_{2}^{a(q-1)})=-y_{a+1}(x_{1},x_{2})=-\delta_{2}(Q_{1,0}^{a+1}).$
Write
$Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})=\sum_{(a,b)}\lambda(a,b)x_{2}^{a(q-1)}x_{3}^{b(q-1)},\quad\lambda(a,b)\in\mathbb{F}_{q},$
then
$\operatorname{Tr}_{B_{3}}^{P(2,1)}\big{(}x_{1}^{q^{m}-1}Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})\big{)}=$ $-\sum_{(a,b)}\lambda(a,b)\delta_{2}(Q_{1,0}^{a+1})x_{3}^{b(q-1)}$
$=$ $-\delta_{2}\big{(}\sum_{(a,b)}\lambda(a,b)x_{1}^{(a+1)(q-1)}x_{2}^{b(q-1)}\big{)}$
$=$ $-\delta_{2}(Q_{1,0}Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})$
$=$ $\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}+1})-Q_{2,0}\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}).$
Finally, we observe that
$\operatorname{Tr}_{P(2,1)}^{G}\big{(}Q_{2,0}\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})\big{)}=\operatorname{Tr}_{P(2,1)}^{G}(Q_{2,0})\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})=0,$
which proves the claim for the fourth family.
Now we consider the fifth family (note that the 6th family already consists of $G$-invariants). A similar argument shows that if $j<[m]_{q}$, then
$\operatorname{Tr}_{P(1,2)}^{G}(x_{1}^{q^{m}-1}\delta_{2}(Q_{1,0}^{j})(x_{2},x_{3}))=-\delta_{2}^{2}(Q_{1,0}^{j+1}).$
Indeed, recall
$\operatorname{Tr}_{B_{3}}^{P(2,1)}(x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3}))=-\sum_{a=j}^{(q^{m}-q)/(q-1)}y_{a+1}(x_{1},x_{2})x_{3}^{q^{m}-q+j(a-1)-a(q-1)}.$
This sum is exactly $-a_{m,3,j+1}=-\delta_{2}^{2}(Q_{1,0}^{j+1})$, which is a $G_{3}$-invariant. The result follows by the transitivity of the transfer. The proof is complete. ∎
The next step is to show that in the third family, one does not need to use all ${\mathcal{D}}_{2}$. The following technical lemma will be used repeatedly.
@@
Let $s,t,i$ be non-negative integers. Then the following equality holds in ${\mathcal{Q}}_{m}(3)$:
@@
Recall that $\delta_{2}(f)$ is a polynomial if $f$ is a polynomial (possibly with more than one variable), which is $G_{1}$-invariant. Hence, the expression under consideration is a polynomial. If $t\geq 1$, unravelling the definition, we see that $\delta_{2}(V_{1}^{s(q-1)}V_{2}^{t(q-1)}\delta_{2}(Q_{1,0}^{i}))$ equals
@@
The first summand is trivial in ${\mathcal{Q}}_{m}(3)$ because the quotient involved is a polynomial. Indeed, in this quotient, the numerator is divisible by $x_{1}$ and vanishes whenever there is a linear relation $x_{2}=a_{1}x_{1}$.
By a similar argument, the second summand is also trivial if $s>1$, and equals $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-1)(q-1)}$ if $s=1$. If $s=0$, it equals
$(t(q-1)-1)x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-2)(q-1)}.$
Now suppose $t=0$, we get
$\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}^{s(q-1)}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{2}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}^{s(q-1)}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{1}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{1},x_{3})}\big{]}\\ =\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{2}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{1}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\big{]}+\\ \frac{x_{1}^{q^{m}}x_{2}^{q^{m}}x_{3}^{i(q-1)}}{L_{2}(x_{1},x_{2})}\Bigg{[}\frac{x_{2}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{2}^{q-1}-x_{3}^{q-1}}-\frac{x_{1}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{1}^{q-1}-x_{3}^{q-1}}\Bigg{]}.$
The first summand is $\delta_{2}^{2}(Q_{1,0}^{s+i})$. The second summand is trivial when $s\leq 1$ and equals $-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(s-2+i)(q-1)}$ if $s\geq 2$. ∎
Corollary 7.3.
If $g\in{\mathcal{D}}_{2}$ and $i\geq 0$, then $\delta_{3}(g\delta_{2}(Q_{1,0}^{i}))$ is a linear combination of polynomials from the following families:
  • $\delta_{2}^{2}(Q_{1,0}^{s})$, $0\leq s<[m]_{q}$.
  • $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{s(q-1)}$, $0\leq s\leq[m]_{q}$.
Proof.
We first observe that by direct inspection, $\delta_{3}(g\delta_{2}(Q_{1,0}^{i}))$ can be written as a quotient of determinants:
$\delta_{3}(g\delta_{2}(Q_{1,0}^{i}))=\frac{\begin{vmatrix}x_{1}^{i(q-1)+1}&x_{2}^{i(q-1)+1}&x_{3}^{i(q-1)+1}\\ x_{1}^{q^{m}}&x_{2}^{q^{m}}&x_{3}^{q^{m}}\\ x_{1}^{q^{m}}g(x_{2},x_{3})&x_{2}^{q^{m}}g(x_{1},x_{3})&x_{3}^{q^{m}}g(x_{1},x_{2})\end{vmatrix}}{L_{3}(x_{1},x_{2},x_{3})}.$
Since
$L_{3}(x_{1},x_{2},x_{3})=$ $V_{3}(x_{1},x_{2},x_{3})L_{2}(x_{1},x_{2})$
the quotient above can be rewritten in the form:
$\varphi(h):=$ $\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{1}^{q^{m}}x_{2}h(x_{1},x_{2},x_{3})\frac{\begin{vmatrix}x_{2}^{i(q-1)+1}&x_{3}^{i(q-1)+1}\\ x_{2}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}h(x_{2},x_{1},x_{3})\frac{\begin{vmatrix}x_{1}^{i(q-1)+1}&x_{3}^{i(q-1)+1}\\ x_{1}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\bigg{]}$
$=$ $\frac{1}{L_{2}(x_{1},x_{2})}(x_{1}^{q^{m}}x_{2}h(x_{1},x_{2},x_{3})\delta_{2}(Q_{1,0}^{i})(x_{2},x_{3})-x_{2}^{q^{m}}x_{1}h(x_{2},x_{1},x_{3})\delta_{2}(Q_{1,0}^{i})(x_{1},x_{3})),$
where
Since $g\in{\mathcal{D}}_{2}$, it can be written in terms of upper triangular invariants
$g=\sum_{s,t}V_{1}^{s(q-1)}V_{2}^{t(q-1)},$
where the sum is taken over certain subset $A\subset\mathbb{N}^{2}$. Then
$h(x_{1},x_{2},x_{3})=\sum_{(s,t)\in A,t\geq 1}x_{2}^{s(q-1)}\bigg{(}\sum_{j=0}^{t-1}V_{2}(x_{2},x_{3})^{(t-1-j)(q-1)}V_{2}(x_{2},x_{1})^{j(q-1)}\bigg{)}.$
Observe that if $j>0$, then both $x_{1}^{q^{m}}x_{2}V_{2}(x_{2},x_{1})^{j(q-1)}/L_{2}$ and $x_{2}^{q^{m}}x_{1}V_{2}(x_{1},x_{2})^{j(q-1)}/L_{2}$ are zero in ${\mathcal{Q}}_{m}$, so the terms in $h$ corresponding to those $j$ do not contribute to $\varphi(h)$. It follows that $\varphi(h)$ is reduced to the sum
$\sum_{(s,t)\in A,t\geq 1}\varphi(x_{2}^{s(q-1)}V_{2}(x_{2},x_{3})^{(t-1)(q-1)}).$
On the other hand, we have an equality which is essentially by definition:
$\varphi(x_{2}^{s(q-1)}V_{2}(x_{2},x_{3})^{(t-1)(q-1)})=\delta_{2}(V_{1}^{s(q-1)}V_{2}^{(t-1)(q-1)}\delta_{2}(Q_{1,0}^{i})).$
We can now use the computation in Lemma 7.2 of the righ-hand side to finish the proof. ∎
We are now ready to replace ${\mathcal{D}}_{2}$ by $\Delta^{m}_{2}$ in the generating set $\mathcal{B}^{\prime}$ of ${\mathcal{Q}}_{m}(3)^{G}$.
@@
The subspace of ${\mathcal{Q}}_{m}(3)^{G}$ spanned by the following three families of elements